Two charges 4q and q are placed at distance l apart. A third charge q is placed at middle of them if resultant force on q is zero then the value of q is

Grade:12

1 Answers

Md hasan
5 years ago
There seems to be a mistake in the question. It`s not for finding Q but for finding the distance of Q from 4q or q.Let the distance from q be x.Apply Coulomb`s law for Electrostatic forces. Q *4q/(l - x)^2 = Q *q/x^2So 4x^2= (l-x)^22x = l - xx = l/3.😊